5=a+4/3 solve for a ... I know the answer but I'm not sure how to set up the equation to show the answer a: 11
Answer by nyc_function(2741)

Any letter by itself is divided by 1.
In other words, a is really a/1. So, a/1 = a because ANYTHING divided by 1 equal ANYTHING. Understand?

This does not change your equation.

The equation 5 = a/1 + 4/3 is EXACTLY the same as 5 = a + 4/3.

Multiply every term on both sides of the equation by 3. This will remove the fraction on the right side.


5 times 3 = a/1 times 3 + (4/3) times 3

15 = 3a + 4

15 - 4 = 3a

11 = 3a

We now divide 11 by 3 to find a.

11/3 = a
The correct answer is not 11 but the improper fraction 11/3.
So, a = 11/3 is the answer.
Understand?